4. Find suitable coordinates for a torus. Are your coordinates orthogonal? Com pute the volume of that torus.

Vector Calculus
Answer #1

The general eq. of Torus. (c-1x2+42)² + ² = at it can bre parametrise as x = (c+ a cosul cosu le sait y = (ct a casu) sinu o